OpenStudy (anonymous):

Factor the trinomial: v^3-6v^2-27v

OpenStudy (anonymous):

v(v-9)(v+3)=0

OpenStudy (anonymous):

v=0,9,-3

OpenStudy (anonymous):

v(v^2 - 6v- 27) = v (v-9)(v+3)